OpenStudy (kmullis6):

WILL FAN AND GIVE MEDAL! Do you mind helping me through this process? I already took the quiz and know the answer, but guessed on it. How was the answer found?

OpenStudy (kmullis6):

The minimum and maximum temperature on a cold day in Lollypop Town can be modeled by the equation below: 2|x − 6| + 14 = 38 What are the minimum and maximum temperatures for this day? x = −9, x = 21 x = −6, x = 18 x = 6, x = −18 No solution --------- The answer is B.

OpenStudy (anonymous):

You need to get |x-6| isolated on one side of the equation. Do you know how?

OpenStudy (kmullis6):

No, but I guess that's why I don't understand how to do this haha (:

OpenStudy (anonymous):

OK lets start. First lets subtract -14 from both sides of the equation. What do you end up with?

OpenStudy (kmullis6):

Like 2 and 38? sorry, new to this

OpenStudy (anonymous):

No prob,,if we subtract 14 from both sides of the equation 2|x-6|+14-14 = 38-14 we end up with 2|x-6| = 24 Do you see?

OpenStudy (kmullis6):

What is the reason for doing that?

OpenStudy (anonymous):

When solving for unknown variables in this case (x) you need to get the unknown on one side of the equation by itself.

OpenStudy (kmullis6):

Ok yea i get it.

OpenStudy (anonymous):

So we have 2|x-6| =24 Now we need to get rid of the 2. Any thoughts?

OpenStudy (kmullis6):

I really have no clue. /:

OpenStudy (anonymous):

Divide both sides by 2...\[\frac{ 2|x-6| }{ 2 } = \frac{ 24 }{ 2 }\]

OpenStudy (anonymous):

o you know what this will leave us witth?D

OpenStudy (kmullis6):

I feel like an idiot because a 5 year old could probably do this lol but what exactly am i dividing? like 24/2 or 2|x-6|/2 or? sorry

OpenStudy (anonymous):

The 2's cancel out on the right side and 24/2 = 12 on the left so we end up with |x-6|=12

OpenStudy (anonymous):

Now to solve for absolute value equation |x-6|=12 we will always have two solutions. The way to solve them is to make your answer both positive and negative like so. x-6 = 12 Solve for x x-6 = -12 Solve for x

OpenStudy (kmullis6):

I get it now, thank you soooo much! :D

OpenStudy (kmullis6):

Actually no, I take it back... I don't get it ): x would be 6 and -18? @nelsonjedi

OpenStudy (kmullis6):

Did I miss something? x is supposed to be -6 and 18.